# Two-dimensional FrBD friction models for rolling contact: extension to linear viscoelasticity

### v2

@@ 2.1 Viscoelastic rheological models
Equations (2) and (12) provide the main physical ingredients required to derive the extended FrBD model, as explained in the following Sect. $2.$
@@ 2.2 Dynamic friction models derivation
Equations (19) and (20) are, however, still finite-dimensional, and cannot accurately capture spatially-varying rolling contact phenomena that evolve over a finite area. The distributed-parameter extension of Eqs. (19) and (20) is worked out next in Sect. 3.
@@ 4.2 Passivity
Passivity arguably constitutes the most important mathematical property of a bona fide friction model, as it ensures consistency with the inherently dissipative nature of the underlying physical phenomenon. For the rolling contact processes described by Eqs. (46) and (47), passivity guarantees that energy is not generated during frictional rolling contact. For the FrBDn+1 models presented in this paper, passivity is also is particularly relevant for control-oriented simulations, as it guarantees stability when they are coupled with mechanical systems or observers. It is worth emphasising that, in this context, passivity should be understood as a property of the process rather than of the friction model itself.
Owing to these preliminary considerations, the corresponding mathematical definition, adopted from [88], is given below.
@@
$-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}\geq\dot{W}\bigl(\bm{u}(\cdot,s)\bigr),\qquad s\in(0,S).$ (55)
@@
$\begin{split}\dot{W}\bigl(\bm{u}(\cdot,s)\bigr)&=-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}+\iint_{\mathscr{C}}p(\bm{x})\bm{z}_{0}^{\mathrm{T}}(\bm{x},s)\bar{\mathbf{K}}_{0}\mathbf{\Psi}\bigl(\bar{\bm{v}}(\bm{x},s)\bigr)\bar{\mathbf{K}}_{0}\bm{z}_{0}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\qquad-\sum_{i=1}^{n}\iint_{\mathscr{C}}p(\bm{x})\bm{z}_{0}^{\mathrm{T}}(\bm{x},s)\bar{\mathbf{K}}_{0}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}_{i}(\bm{x},s)}{\operatorname{d\!}{s}}}\operatorname{d\!}\bm{x}\\ &\qquad+\sum_{i=1}^{n}\iint_{\mathscr{C}}p(\bm{x})\biggl(\bar{\mathbf{K}}_{0}\bm{z}_{0}-V_{\textnormal{r}}\bar{\mathbf{C}}_{i}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}_{i}(\bm{x},s)}{\operatorname{d\!}{s}}}\biggr)^{\mathrm{T}}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}_{i}(\bm{x},s)}{\operatorname{d\!}{s}}}\operatorname{d\!}\bm{x}\\ &\qquad-\dfrac{1}{2}\sum_{i=0}^{n}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}_{i}^{\mathrm{T}}(\bm{x},s)\bar{\mathbf{K}}_{i}\bm{z}_{i}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (58)
@@
$\begin{split}\dot{W}\bigl(\bm{u}(\cdot,s)\bigr)&=-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}+\iint_{\mathscr{C}}p(\bm{x})\bm{z}_{0}^{\mathrm{T}}(\bm{x},s)\bar{\mathbf{K}}_{0}\mathbf{\Psi}\bigl(\bar{\bm{v}}(\bm{x},s)\bigr)\bar{\mathbf{K}}_{0}\bm{z}_{0}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\qquad-\sum_{i=1}^{n}\iint_{\mathscr{C}}p(\bm{x})V_{\textnormal{r}}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}_{i}^{\mathrm{T}}(\bm{x},s)}{\operatorname{d\!}{s}}}\bar{\mathbf{C}}_{i}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}_{i}(\bm{x},s)}{\operatorname{d\!}{s}}}\operatorname{d\!}\bm{x}\\ &\qquad+\dfrac{1}{2}\sum_{i=0}^{n}\iint_{\mathscr{C}}\bigl(\nabla_{\bm{x}}\cdot p(\bm{x})\bar{\bm{V}}(\bm{x})\bigr)\bm{z}_{i}^{\mathrm{T}}(\bm{x},s)\bar{\mathbf{K}}_{i}\bm{z}_{i}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (59)
@@
If $p(\bm{x})\geq p_{\textnormal{min}}$ and $\nabla_{\bm{x}}\cdot p(\bm{x})\bar{\bm{V}}(\bm{x})\leq-\gamma$, with $p_{\textnormal{min}},\gamma\in\mathbb{R}_{>0}$, for all $\bm{x}\in\mathscr{C}$, then the storage functions $W(\bm{u}(\cdot,s))$ in Eqs. (51) and (56) are indeed a Lyapunov functions, and ensures ISS for the PDEs (46) and (47). In this context, it is worth observing that, for $n=0$, the storage functions (51) and (56) reduce to that obtained for the FrBD1-KV model introduced in [88].
@@ 6 Conclusions
  • [5] Kalker $J$. Three-Dimensional Elastic Bodies in Rolling Contact. 1st ed. Springer, Dordrecht (1990). Available from: https://doi.org/10.1007/978-94-015-7889-9.
  • [6] Kalker $J$. Rolling Contact Phenomena. In: Jacobson B, Kalker $J$ (eds) Rolling Contact Phenomena. International Centre for Mechanical Sciences, vol. 411. Springer, Vienna (2000).
@@
  • [32] Kalker $J$. Transient phenomena in two elastic cylinders rolling over each other with dry friction. J Appl Mech 37(3):677–688 (1970). Available from: https://doi.org/10.1115/1.3408597.
  • [33] Kalker $J$. Transient rolling contact phenomena. ASLE Trans 14(3):177–184 (1971). Available from: https://doi.org/10.1080/05698197108983240.
  • [34] Kalker $J$. On the Rolling Contact of Two Elastic Bodies in the Presence of Dry Friction. PhD thesis, TH Delft (1967).
@@
  • [37] Panek C, Kalker $J$. Three-dimensional Contact of a Rigid Roller Traversing a Viscoelastic Half Space. J. Inst. Maths Applies 26:299-313 (1980).
  • [38] Kalker $J$. Viscoelastic Multilayered Cylinders Rolling With Dry Friction. ASME. J. Appl. Mech. September 58(3):666–679 (1991). Available from: https://doi.org/10.1115/1.2897247
@@
  • [51] Goryacheva IG, Gubenko $M$, Makhovskaya $Y$. Sliding of a spherical indenter on a viscoelastic foundation with the forces of molecular attraction taken into account. Prikl. J. Appl. Mech. Tech. Phys., 55(1):81–88 (2014).
@@
  • [55] Kalker $J$. Simplified theory of rolling contact. Delft Progress Report, Series C (1973), 1–10.
  • [56] Kalker $J$. Railway wheel and automotive tyre. Veh Syst Dyn 5(15):255–269 (1979).

### v3

@@ 2.1 Viscoelastic rheological models
Equations (2) and (12) provide the main physical ingredients required to derive the extended FrBD model, as explained in the following Sect. 2.2.
@@ 2.2 Dynamic friction models derivation
A second interesting consideration pertains the dissipative nature of the model, which should be evaulated considering the true sliding velocity $\bm{v}_{\textnormal{s}}(\dot{\bm{z}},\bm{v}_{\textnormal{r}})$. In particular, Eq. (18) yields
$\bm{v}_{\textnormal{s}}\bigl(\dot{\bm{z}}(t),\bm{v}_{\textnormal{r}}(t)\bigr)=\bm{v}_{\textnormal{r}}(t)+\dot{\bm{z}}(t)=-\mathbf{M}^{-2}\bigl(\bm{v}_{\textnormal{r}}(t)\bigr)\mathinner{\!\left\lVert\mathbf{M}\bigl(\bm{v}_{\textnormal{r}}(t)\bigr)\bm{v}_{\textnormal{r}}(t)\right\rVert}_{2,\varepsilon}\bm{f}(t),\qquad t\in[0,T].$ (24)
Consequently, for $p\in C^{0}([0,T];\in\mathbb{R}_{\geq 0})$, the dissipated power reads
$-p(t)\bm{f}^{\mathrm{T}}(t)\bm{v}_{\textnormal{s}}\bigl(\dot{\bm{z}}(t),\bm{v}_{\textnormal{r}}(t)\bigr)=p(t)\bm{f}^{\mathrm{T}}(t)\mathbf{M}^{-2}\bigl(\bm{v}_{\textnormal{r}}(t)\bigr)\mathinner{\!\left\lVert\mathbf{M}\bigl(\bm{v}_{\textnormal{r}}(t)\bigr)\bm{v}_{\textnormal{r}}(t)\right\rVert}_{2,\varepsilon}\bm{f}(t)\geq 0,\qquad t\in[0,T].$ (25)
Equations (19) and (20) appear therefore to preserve the natural properties of the friction force (13). They are, however, still finite-dimensional, and cannot accurately capture spatially-varying rolling contact phenomena that evolve over a finite area. The distributed-parameter extension of Eqs. (19) and (20) is worked out next in Sect. 3.
@@ 4.2 Passivity
Dissipativity and passivity arguably constitute the most important mathematical properties of a bona fide friction model, as they ensure consistency with the inherently dissipative nature of the underlying physical phenomenon. For the rolling contact processes described by Eqs. (48) and (49), dissipativity expresses the fact that friction leads to a net energy loss within the contact region, whereas passivity ensures that no energy is generated during frictional rolling contact. For the FrBDn+1 models presented in this paper, passivity is also particularly relevant for control-oriented simulations, as it guarantees stability when coupled with mechanical systems or observers. It is worth emphasising that, in this context, passivity should be understood as a property of the process rather than of the friction model itself. Starting with dissipativity, observing that
$\bm{v}_{\textnormal{s}}(\bm{x},s)=V_{\textnormal{r}}(s)\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}(\bm{x},s)}{\operatorname{d\!}{s}}}+V_{\textnormal{r}}(s)\bar{\bm{v}}_{\textnormal{r}}(\bm{x},s),\qquad s\in[0,S],$ (51)
and in analogy to what done in Sect. 2.2, for $p\in C^{0}(\mathscr{C}\times[0,S];\mathbb{R}_{\geq 0})$ and $V_{\textnormal{r}}\in C^{0}([0,S];[V_{\textnormal{min}},V_{\textnormal{max}}])$, the dissipated power may be inferred to be
$-\iint_{\mathscr{C}}p(\bm{x},s)\bm{f}^{\mathrm{T}}(\bm{x},s)\bm{v}_{\textnormal{s}}(\bm{x},s)\operatorname{d\!}\bm{x}=-\iint_{\mathscr{C}}V_{\textnormal{r}}(s)p(\bm{x},s)\bm{f}^{\mathrm{T}}(\bm{x},s)\mathbf{\Psi}\bigl(\bar{\bm{v}}_{\textnormal{r}}(\bm{x},s),s\bigr)\bm{f}(\bm{x},s)\operatorname{d\!}\bm{x}\geq 0,\qquad s\in[0,S],$ (52)
which is consistent with the nature of the friction process. Equation (52) evaluates the dissipation rate considering the physically correct velocity, namely the sliding velocity of the tip of the bristles. However, from a dynamical system theory perspective, the input-output behaviour of the PDEs (48) and (49) may also be studied regarding the velocity $\bm{v}(\bm{x},s)$ (or equivalently $\bar{\bm{v}}(\bm{x},s)$) as the input. Owing to these preliminary considerations, the mathematical definitions of dissipativity and passivity, adopted from [88], are given below.
@@
$-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}\geq\mathinner{\dfrac{\operatorname{d\!}{}W\bigl(\bm{u}(\cdot,s)\bigr)}{\operatorname{d\!}{s}}},\qquad s\in(0,S).$ (59)
@@
$\begin{split}\mathinner{\dfrac{\operatorname{d\!}{}W\bigl(\bm{u}(\cdot,s)\bigr)}{\operatorname{d\!}{s}}}&=-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}+\iint_{\mathscr{C}}p(\bm{x})\bm{z}_{0}^{\mathrm{T}}(\bm{x},s)\bar{\mathbf{K}}_{0}\mathbf{\Psi}\bigl(\bar{\bm{v}}(\bm{x},s)\bigr)\bar{\mathbf{K}}_{0}\bm{z}_{0}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\qquad-\sum_{i=1}^{n}\iint_{\mathscr{C}}p(\bm{x})\bm{z}_{0}^{\mathrm{T}}(\bm{x},s)\bar{\mathbf{K}}_{0}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}_{i}(\bm{x},s)}{\operatorname{d\!}{s}}}\operatorname{d\!}\bm{x}\\ &\qquad+\sum_{i=1}^{n}\iint_{\mathscr{C}}p(\bm{x})\biggl(\bar{\mathbf{K}}_{0}\bm{z}_{0}-V_{\textnormal{r}}\bar{\mathbf{C}}_{i}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}_{i}(\bm{x},s)}{\operatorname{d\!}{s}}}\biggr)^{\mathrm{T}}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}_{i}(\bm{x},s)}{\operatorname{d\!}{s}}}\operatorname{d\!}\bm{x}\\ &\qquad-\dfrac{1}{2}\sum_{i=0}^{n}\iint_{\mathscr{C}}p(\bm{x})\bar{\bm{V}}(\bm{x})\cdot\nabla_{\bm{x}}\Bigl(\bm{z}_{i}^{\mathrm{T}}(\bm{x},s)\bar{\mathbf{K}}_{i}\bm{z}_{i}(\bm{x},s)\Bigr)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (62)
@@
$\begin{split}\mathinner{\dfrac{\operatorname{d\!}{}W\bigl(\bm{u}(\cdot,s)\bigr)}{\operatorname{d\!}{s}}}&=-\bigl\langle p(\cdot)\bm{f}(\cdot,s),\bar{\bm{v}}(\cdot,s)\bigr\rangle_{L^{2}(\mathring{\mathscr{C}};\mathbb{R}^{2})}+\iint_{\mathscr{C}}p(\bm{x})\bm{z}_{0}^{\mathrm{T}}(\bm{x},s)\bar{\mathbf{K}}_{0}\mathbf{\Psi}\bigl(\bar{\bm{v}}(\bm{x},s)\bigr)\bar{\mathbf{K}}_{0}\bm{z}_{0}(\bm{x},s)\operatorname{d\!}\bm{x}\\ &\qquad-\sum_{i=1}^{n}\iint_{\mathscr{C}}p(\bm{x})V_{\textnormal{r}}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}_{i}^{\mathrm{T}}(\bm{x},s)}{\operatorname{d\!}{s}}}\bar{\mathbf{C}}_{i}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}_{i}(\bm{x},s)}{\operatorname{d\!}{s}}}\operatorname{d\!}\bm{x}\\ &\qquad+\dfrac{1}{2}\sum_{i=0}^{n}\iint_{\mathscr{C}}\bigl(\nabla_{\bm{x}}\cdot p(\bm{x})\bar{\bm{V}}(\bm{x})\bigr)\bm{z}_{i}^{\mathrm{T}}(\bm{x},s)\bar{\mathbf{K}}_{i}\bm{z}_{i}(\bm{x},s)\operatorname{d\!}\bm{x},\qquad s\in(0,S).\end{split}$ (63)
@@
If $p(\bm{x})\geq p_{\textnormal{min}}$ and $\nabla_{\bm{x}}\cdot p(\bm{x})\bar{\bm{V}}(\bm{x})\leq-\gamma$, with $p_{\textnormal{min}},\gamma\in\mathbb{R}_{>0}$, for all $\bm{x}\in\mathscr{C}$, then the storage functions $W(\bm{u}(\cdot,s))$ in Eqs. (55) and (60) are indeed a Lyapunov functions, and ensure ISS for the PDEs (48) and (49). In this context, it is worth observing that, for $n=0$, the storage functions (55) and (60) reduce to that obtained for the FrBD1-KV model introduced in [88]. In fact, under the condition (54), passivity properties for the FrBD1-KV model, for which
$\bm{f}(\bm{x},s)=\bar{\mathbf{K}}\bm{z}(\bm{x},s)+V_{\textnormal{r}}(s)\bar{\mathbf{C}}\mathinner{\dfrac{\operatorname{d\!}{}\bm{z}(\bm{x},s)}{\operatorname{d\!}{s}}},$ (64)
may be inferred directly from (52) for all $\mathbf{Sym}_{2}(\mathbb{R})\ni\bar{\mathbf{K}}\succ\mathbf{0}$, $\mathbf{Sym}_{2}(\mathbb{R})\ni\bar{\mathbf{C}}\succeq\mathbf{0}$, and $V_{\textnormal{r}}\in[V_{\textnormal{min}},V_{\textnormal{max}}]$, generalising the findings of [88].
@@ 6 Conclusions
  • [5] Kalker JJ. Three-Dimensional Elastic Bodies in Rolling Contact. 1st ed. Springer, Dordrecht (1990). Available from: https://doi.org/10.1007/978-94-015-7889-9.
  • [6] Kalker JJ. Rolling Contact Phenomena. In: Jacobson B, Kalker JJ (eds) Rolling Contact Phenomena. International Centre for Mechanical Sciences, vol. 411. Springer, Vienna (2000).
@@
  • [32] Kalker JJ. Transient phenomena in two elastic cylinders rolling over each other with dry friction. J Appl Mech 37(3):677–688 (1970). Available from: https://doi.org/10.1115/1.3408597.
  • [33] Kalker JJ. Transient rolling contact phenomena. ASLE Trans 14(3):177–184 (1971). Available from: https://doi.org/10.1080/05698197108983240.
  • [34] Kalker JJ. On the Rolling Contact of Two Elastic Bodies in the Presence of Dry Friction. PhD thesis, TH Delft (1967).
@@
  • [37] Panek C, Kalker JJ. Three-dimensional Contact of a Rigid Roller Traversing a Viscoelastic Half Space. J. Inst. Maths Applies 26:299-313 (1980).
  • [38] Kalker JJ. Viscoelastic Multilayered Cylinders Rolling With Dry Friction. ASME. J. Appl. Mech. September 58(3):666–679 (1991). Available from: https://doi.org/10.1115/1.2897247
@@
  • [51] Goryacheva IG, Gubenko MM, Makhovskaya YY. Sliding of a spherical indenter on a viscoelastic foundation with the forces of molecular attraction taken into account. Prikl. J. Appl. Mech. Tech. Phys., 55(1):81–88 (2014).
@@
  • [55] Kalker JJ. Simplified theory of rolling contact. Delft Progress Report, Series C (1973), 1–10.
  • [56] Kalker JJ. Railway wheel and automotive tyre. Veh Syst Dyn 5(15):255–269 (1979).